# Explainability in Neural Networks for Natural Language Processing Tasks

### v1

@@ 1 Introduction
In today’s technology-driven world, natural language processing (NLP) has become a cornerstone of many applications, enabling systems such as voice assistants, email filters, chatbots, and predictive text to facilitate seamless communication [Young et al., 2018]. These advances, powered by neural network models, have revolutionized how machines understand and process human language. However, despite their remarkable success, neural network models are often considered opaque black-box systems, leaving users and researchers with limited insights into how these models generate their predictions. This lack of transparency raises critical concerns, particularly in high-stakes domains such as healthcare, finance, and law, where accountability and trust are paramount. The concept of interpretability—defined as a model’s ability to provide human-understandable reasoning for its predictions [Lipton, 2018]—has emerged as a vital area of research. Understanding the inner workings of neural network models not only fosters trust but also aids in diagnosing errors, identifying biases, and ensuring ethical AI deployment. However, achieving interpretability for neural network-based NLP models is especially challenging due to the complexity of their layered architectures and the intricacies of language itself [Alishahi et al., 2019]. To address these challenges, this study employs Local Interpretable Model-Agnostic Explanations (LIME), a widely used technique for explaining black-box models. LIME generates localized explanations by approximating a model’s behavior for individual instances, focusing on the contributions of specific features to predictions. This approach is particularly valuable for text classification tasks, as it allows for a granular understanding of how words influence a model’s output. While LIME provides significant insights at the instance level, it fails to address global interpretability and capture interactions between features. This study explores the application of LIME to a multi-layer perceptron (MLP) neural network trained on text classification, highlighting the technique’s strengths and limitations. By shedding light on the decision-making process of neural NLP models, this study contributes to the broader goal of making AI models more transparent, trustworthy, and accountable.
@@ 2.1 Why Interpretability Matters
In machine learning, model interpretability is a cornerstone for understanding how systems operate, particularly in real-world applications that influence critical decisions [Mersha et al., 2024a]. For example, in fields such as healthcare, finance, and legal systems, the impact of model predictions can be profound. However, traditional evaluation metrics like classification accuracy often fail to provide meaningful insights into a model’s behavior and decision-making process [Doshi-Velez and Kim, 2017]. Interpretability helps ensure that models align with human expectations and plays a crucial role in identifying biases, uncovering vulnerabilities, and debugging unexpected outcomes. By making models more transparent and understandable, interpretability fosters trust, enhances accountability, and enables more informed decision-making.
@@ 2.3 Local Interpretable Model-agnostic Explanation
Local Interpretable Model-Agnostic Explanations (LIME) is a widely recognized technique for explaining black-box models by approximating their behavior locally around individual predictions [Yigezu et al., 2024]. Instead of explaining the entire model globally, LIME focuses on generating localized, interpretable representations that make sense to humans. Its functionality varies depending on the type of data it is applied to. For text data, LIME creates explanations by selectively removing or masking words and analyzing their impact on predictions. For tabular data, it generates perturbations by altering feature values, while for images, it uses superpixel segmentation to modify specific regions. By training local surrogate models to mimic the behavior of the original black-box model, LIME provides instance-specific insights into how individual features contribute to predictions. Despite its advantages in localized interpretability, LIME’s limitations include its inability to address global model behavior and its dependency on the quality of surrogate models.
@@
Figure 1: LIME explanation logic [Ribeiro et al., 2016]
@@ 3 Literature Review
Neural networks with multiple layers can automatically learn complex patterns and representations from large datasets, with applications in computer vision, natural language processing (NLP), healthcare, and autonomous systems [AlShami et al., 2024]. Recent research has extensively focused on the interpretability and analysis of neural network models for natural language processing (NLP), addressing the critical challenge of understanding their complex decision-making processes. Interpretability has emerged as a cornerstone of explainable AI (XAI), particularly for NLP models deployed in high-stakes domains.
Lipton [Lipton, 2018] conducted a seminal study that clarified the concept of interpretability in machine learning and underscored its significance for neural NLP models. The study established a foundational framework for defining interpretability, including taxonomies and evaluation criteria, highlighting its pivotal role in fostering trust and accountability in AI systems.
Ribeiro et al. [Ribeiro et al., 2016] and Montavon et al. [Montavon et al., 2018] explored a range of neural network interpretation techniques, including Local Interpretable Model-Agnostic Explanations (LIME), Gradient-Based Sensitivity Analysis, and SHapley Additive exPlanations (SHAP). Ribeiro et al. emphasized LIME’s versatility across data modalities, while Montavon et al. advanced the understanding of gradient-based methods by leveraging their compatibility with deep neural network architectures.
Belinkov et al. [Belinkov et al., 2020] provided a comprehensive discussion on neural network interpretability, covering structural analyses through probing classifiers, behavioral evaluations via test suites, and interactive visualizations for debugging model behavior. Their work critically analyzed the strengths and weaknesses of various interpretability methods, offering insights into their applicability and limitations, such as the inability to capture contextual dependencies in language tasks fully.
@@ 4 Methodology
The agnostic model explanation method is the most common explanation technique that can be applied to different types of ML models [Gholizadeh and Zhou, 2021]. Local Interpretable Model-agnostic Explanations (LIME) is an agnostic method [Ribeiro et al., 2016]. LIME focuses on training local surrogate models to explain distinct predictions rather than training a global surrogate model.
@@ 5.1 Text Classification
Text classification, also known as text categorization, involves organizing textual data into predefined categories [Shah et al., 2019]. This task is foundational in natural language processing (NLP) and finds applications in spam detection, sentiment analysis, topic labeling, and more. For this study, a Multilayer Perceptron (MLP) was employed as the classifier due to its ability to learn complex patterns in textual data through non-linear transformations. The MLP architecture is particularly suited for classification tasks, offering flexibility and robustness when combined with effective preprocessing and feature engineering techniques.
@@ 5.1.1 Dataset
For this study, 80% of the data (16,000 documents) is used for training, and 20% (4,000 documents) for testing. Challenges: Overlapping vocabularies between categories and noisy metadata make classification challenging but realistic. This dataset’s diversity and complexity make it ideal for evaluating the classification model and testing LIME-based explainability [Tonja et al., 2023].
@@ 5.1.2 Prepossessing
Preprocessing is a critical step in preparing textual data for analysis, ensuring consistency and reducing noise to improve model performance. The process began with case normalization, converting all text to lowercase to eliminate case sensitivity and reduce feature dimensionality. Next, excess whitespace between words, along with punctuation marks and special characters, was removed to focus on meaningful content. Stopword removal was applied to filter out frequently used but semantically insignificant words, such as ”the,” ”is,” and ”and,” using a predefined stopword list. The text was then tokenized into word-level tokens to prepare for feature extraction. To further enhance consistency, stemming and lemmatization were applied to reduce words to their root forms, with lemmatization prioritized for its contextual accuracy. Additionally, irrelevant metadata, such as email headers, signatures, and quoted text, was stripped from the documents to eliminate noise. Finally, text length filtering was performed to remove extremely short or excessively long documents, ensuring uniformity across the dataset. These preprocessing techniques collectively transformed the raw text into a structured format suitable for feature extraction and classification, optimizing the dataset for downstream tasks.
@@ 5.2 LIME Model Explanation
The LIME model explanations use an interpretable data representation, not the actual features used by the classifier model, because the interpretable data representation is human-understandable. For example, a binary vector (bag of words) is an interpretable representation of the presence or absence of a word [Ribeiro et al., 2016].
The LIME output is a list explanation that tells the effect of each feature on the prediction of a given instance. This is the notion of local interpretability. Based on this local interpretability, we can also determine which feature has a high contribution to changing the prediction output. The LIME explanation is discussed as follows. We select an instance (document) by its index and feed it to the classifier model. The neural networks classifier predicts the probability of a selected instance (document index 20) class being atheism is 0.42, and class Christian is 0.58 as shown from experiment 1. Now the question comes to a neural network, “Why should I trust you?”.
@@ 5.2.1 Train the LIME Model
5.2.2 Validate the LIME Model
The validation dataset is preprocessed with special attention because the 20newsgroup dataset is problematic in validating the model [Ribeiro et al., 2016]. This preprocessing implementation is one of the contributions that enabled us to use the same dataset for training and validation.
@@ 6 Future Work
In this project, we utilized the Bag-of-Words (BoW) representation to explain the MLP classifier model. However, future work will focus on leveraging more sophisticated models to enhance explainability and performance. Specifically, adopting transformer-based models like BERT (Bidirectional Encoder Representations from Transformers) or Sentence Transformers can provide deeper contextual understanding and more accurate feature representations [Mersha et al., 2024b].

### v2

@@ Abstract
Keywords: Explainable AI, Neural Networks, Natural Language Processing, interpretability, black-box models, explainability techniques, LIME, explainable machine learning
@@ 1 Introduction
In today’s technology-driven world, natural language processing (NLP) has become a cornerstone of many applications, enabling systems such as voice assistants, email filters, chatbots, and predictive text to facilitate seamless communication Young et al. (2018). These advances, powered by neural network models, have revolutionized how machines understand and process human language. However, despite their remarkable success, neural network models are often considered opaque black-box systems, leaving users and researchers with limited insights into how these models generate their predictions. This lack of transparency raises critical concerns, particularly in high-stakes domains such as healthcare, finance, and law, where accountability and trust are paramount. The concept of interpretability—defined as a model’s ability to provide human-understandable reasoning for its predictions Lipton (2018)—has emerged as a vital area of research. Understanding the inner workings of neural network models not only fosters trust but also aids in diagnosing errors, identifying biases, and ensuring ethical AI deployment. However, achieving interpretability for neural network-based NLP models is especially challenging due to the complexity of their layered architectures and the intricacies of language itself Alishahi et al. (2019). To address these challenges, this study employs Local Interpretable Model-Agnostic Explanations (LIME), a widely used technique for explaining black-box models. LIME generates localized explanations by approximating a model’s behavior for individual instances, focusing on the contributions of specific features to predictions. This approach is particularly valuable for text classification tasks, as it allows for a granular understanding of how words influence a model’s output. While LIME provides significant insights at the instance level, it fails to address global interpretability and capture interactions between features. This study explores the application of LIME to a multi-layer perceptron (MLP) neural network trained on text classification, highlighting the technique’s strengths and limitations. By shedding light on the decision-making process of neural NLP models, this study contributes to the broader goal of making AI models more transparent, trustworthy, and accountable.
@@ 2.1 Why Interpretability Matters
In machine learning, model interpretability is a cornerstone for understanding how systems operate, particularly in real-world applications that influence critical decisions Mersha et al. (2024a). For example, in fields such as healthcare, finance, and legal systems, the impact of model predictions can be profound. However, traditional evaluation metrics like classification accuracy often fail to provide meaningful insights into a model’s behavior and decision-making process Doshi-Velez and Kim (2017). Interpretability helps ensure that models align with human expectations and plays a crucial role in identifying biases, uncovering vulnerabilities, and debugging unexpected outcomes. By making models more transparent and understandable, interpretability fosters trust, enhances accountability, and enables more informed decision-making.
@@ 2.3 Local Interpretable Model-agnostic Explanation
Local Interpretable Model-Agnostic Explanations (LIME) is a widely recognized technique for explaining black-box models by approximating their behavior locally around individual predictions Yigezu et al. (2024). Instead of explaining the entire model globally, LIME focuses on generating localized, interpretable representations that make sense to humans. Its functionality varies depending on the type of data it is applied to. For text data, LIME creates explanations by selectively removing or masking words and analyzing their impact on predictions. For tabular data, it generates perturbations by altering feature values, while for images, it uses superpixel segmentation to modify specific regions. By training local surrogate models to mimic the behavior of the original black-box model, LIME provides instance-specific insights into how individual features contribute to predictions. Despite its advantages in localized interpretability, LIME’s limitations include its inability to address global model behavior and its dependency on the quality of surrogate models.
@@
Figure 1: LIME explanation logic Ribeiro et al. (2016)
@@ 3 Literature Review
Neural networks with multiple layers can automatically learn complex patterns and representations from large datasets, with applications in computer vision, natural language processing (NLP), healthcare, and autonomous systems AlShami et al. (2024). Recent research has extensively focused on the interpretability and analysis of neural network models for natural language processing (NLP), addressing the critical challenge of understanding their complex decision-making processes. Interpretability has emerged as a cornerstone of explainable AI (XAI), particularly for NLP models deployed in high-stakes domains.
Lipton Lipton (2018) conducted a seminal study that clarified the concept of interpretability in machine learning and underscored its significance for neural NLP models. The study established a foundational framework for defining interpretability, including taxonomies and evaluation criteria, highlighting its pivotal role in fostering trust and accountability in AI systems.
Ribeiro et al. Ribeiro et al. (2016) and Montavon et al. Montavon et al. (2018) explored a range of neural network interpretation techniques, including Local Interpretable Model-Agnostic Explanations (LIME), Gradient-Based Sensitivity Analysis, and SHapley Additive exPlanations (SHAP). Ribeiro et al. emphasized LIME’s versatility across data modalities, while Montavon et al. advanced the understanding of gradient-based methods by leveraging their compatibility with deep neural network architectures.
Belinkov et al. Belinkov et al. (2020) provided a comprehensive discussion on neural network interpretability, covering structural analyses through probing classifiers, behavioral evaluations via test suites, and interactive visualizations for debugging model behavior. Their work critically analyzed the strengths and weaknesses of various interpretability methods, offering insights into their applicability and limitations, such as the inability to capture contextual dependencies in language tasks fully.
@@ 4 Methodology
The agnostic model explanation method is the most common explanation technique that can be applied to different types of ML models Gholizadeh and Zhou (2021). Local Interpretable Model-agnostic Explanations (LIME) is an agnostic method Ribeiro et al. (2016). LIME focuses on training local surrogate models to explain distinct predictions rather than training a global surrogate model.
@@ 5.1 Text Classification
Text classification, also known as text categorization, involves organizing textual data into predefined categories Shah et al. (2019). This task is foundational in natural language processing (NLP) and finds applications in spam detection, sentiment analysis, topic labeling, and more. For this study, a Multilayer Perceptron (MLP) was employed as the classifier due to its ability to learn complex patterns in textual data through non-linear transformations. The MLP architecture is particularly suited for classification tasks, offering flexibility and robustness when combined with effective preprocessing and feature engineering techniques.
@@ 5.1.1 Dataset
For this study, 80% of the data (16,000 documents) is used for training, and 20% (4,000 documents) for testing. Challenges: Overlapping vocabularies between categories and noisy metadata make classification challenging but realistic. This dataset’s diversity and complexity make it ideal for evaluating the classification model and testing LIME-based explainability Tonja et al. (2023).
@@ 5.1.2 Prepossessing
Preprocessing is a critical step in preparing textual data for analysis, ensuring consistency and reducing noise to improve model performance. The process began with case normalization, converting all text to lowercase to eliminate case sensitivity and reduce feature dimensionality. Next, excess whitespace between words, along with punctuation marks and special characters, was removed to focus on meaningful content. Stopword removal was applied to filter out frequently used but semantically insignificant words, such as "the," "is," and "and," using a predefined stopword list. The text was then tokenized into word-level tokens to prepare for feature extraction. To further enhance consistency, stemming and lemmatization were applied to reduce words to their root forms, with lemmatization prioritized for its contextual accuracy. Additionally, irrelevant metadata, such as email headers, signatures, and quoted text, was stripped from the documents to eliminate noise. Finally, text length filtering was performed to remove extremely short or excessively long documents, ensuring uniformity across the dataset. These preprocessing techniques collectively transformed the raw text into a structured format suitable for feature extraction and classification, optimizing the dataset for downstream tasks.
@@ 5.2 LIME Model Explanation
The LIME model explanations use an interpretable data representation, not the actual features used by the classifier model, because the interpretable data representation is human-understandable. For example, a binary vector (bag of words) is an interpretable representation of the presence or absence of a word Ribeiro et al. (2016).
The LIME output is a list explanation that tells the effect of each feature on the prediction of a given instance. This is the notion of local interpretability. Based on this local interpretability, we can also determine which feature has a high contribution to changing the prediction output. The LIME explanation is discussed as follows. We select an instance (document) by its index and feed it to the classifier model. The neural networks classifier predicts the probability of a selected instance (document index 20) class being atheism is 0.42, and class Christian is 0.58 as shown from experiment 1. Now the question comes to a neural network, “Why should I trust you?".
@@
The validation dataset is preprocessed with special attention because the 20newsgroup dataset is problematic in validating the model Ribeiro et al. (2016). This preprocessing implementation is one of the contributions that enabled us to use the same dataset for training and validation.
@@ 6 Future Work
In this project, we utilized the Bag-of-Words (BoW) representation to explain the MLP classifier model. However, future work will focus on leveraging more sophisticated models to enhance explainability and performance. Specifically, adopting transformer-based models like BERT (Bidirectional Encoder Representations from Transformers) or Sentence Transformers can provide deeper contextual understanding and more accurate feature representations Mersha et al. (2024b).